OpenStudy (anonymous):

Write as an equivalent expression using the properties of logs: logb(³√5x/12y^4) *the b is little

OpenStudy (anonymous):

\[\log_b(\sqrt[3]{\frac{5x}{12y^4}})\] can be rewritten as \[\log_b((\frac{5x}{12y^4})^{1/3})\] In a logarithmic function, you can move the exponent to the coefficient. \[\frac{1}{3}\log_b(\frac{5x}{12y^4})\] Division in a logarthimic function is seen as subtraction outside. (I'll provide an example) \[log_b\frac{y^3}{x^2}=log_b(y^3)-log_b(x^2)\] Once you do that, you can once again move the exponent to the coefficient. (continuing with that example, here's the finished product) \[log_b(y^3)-log_b(x^2)=3log_b(y)-2log_b(x)\]